# Weight Space Representation Learning on Diverse NeRF Architectures

### v2

@@ Abstract
Neural Radiance Fields (NeRFs) have emerged as a groundbreaking paradigm for representing 3D objects and scenes by encoding shape and appearance information into the weights of a neural network. Recent studies have demonstrated that these weights can be used as input for frameworks designed to address deep learning tasks; however, such frameworks require NeRFs to adhere to a specific, predefined architecture. In this paper, we introduce the first framework capable of processing NeRFs with diverse architectures and performing inference on architectures unseen at training time. We achieve this by training a Graph Meta-Network within an unsupervised representation learning framework, and show that a contrastive objective is conducive to obtaining an architecture-agnostic latent space. In experiments conducted across 13 NeRF architectures belonging to three families (MLPs, tri-planes, and, for the first time, hash tables), our approach demonstrates robust performance in classification and retrieval tasks involving multiple architectures, even unseen at training time, while also exceeding the results of existing frameworks limited to single architectures.
@@ 1 Introduction
Neural Radiance Fields (NeRFs) [26] have emerged over the last few years as a new paradigm for representing 3D objects and scenes [36]. A NeRF is a neural network trained on a collection of images to map 3D coordinates to color and density values, which can then be used to synthesize novel views of the underlying object or scene via volume rendering. Due to their continuous nature, NeRFs can encode an arbitrary number of images at any resolution into a finite number of neural network weights, thus decoupling the number of observations and their spatial resolution from the memory required to store the 3D representation. As a result, NeRFs hold the potential to become a standard tool for storing and communicating 3D information, as supported by the recent publication of several NeRF datasets [12, 18, 38, 7].
@@
With the rise of NeRFs as a new data format, whether and how it is possible to perform traditional deep learning tasks on them has become an increasingly relevant research question. The naive solution to this problem involves rendering views of the underlying object from its NeRF representation and leveraging existing neural architectures designed to process images. However, this procedure requires additional computation time and several decisions that are likely to impact its outcome, such as the number of views to render, their viewpoint, and their resolution. A more elegant and efficient approach explored by recent works relies instead on performing tasks on NeRFs by processing their weights as input, therefore requiring no rendering step. This strategy is adopted by nf2vec [38], a representation learning framework that learns to map NeRF weights to latent vectors by minimizing a rendering loss, where NeRFs are standard Multi-Layer Perceptrons (MLPs). These vectors are then used as input to deep learning pipelines for downstream tasks. A related approach is proposed by Cardace et al. [7], who, instead of employing traditional MLPs, leverage tri-planar NeRFs [8], where input coordinates are projected onto three orthogonal planes of learnable features to compute the input for an MLP.
@@
Both nf2vec [38] and the method by Cardace et al. [7], however, are designed to ingest a specific type of NeRF architecture (i.e. MLPs with fixed hidden dimensions in nf2vec and tri-planes with fixed spatial resolution in Cardace et al.), and are thus unable to process diverse input architectures. In the context of a research domain where new NeRF designs are constantly being explored [36], this assumption strongly limits their applicability. The issue of handling arbitrary architectures has recently been studied in the broader research field on meta-networks, i.e. neural networks that process other neural networks as input. Specifically, Graph Meta-Networks (GMNs) have been proposed [23, 20], namely Graph Neural Networks (GNNs) that can ingest any neural architecture as long as it can be first converted into a graph. Yet, these works do not experiment with NeRFs as input to GMNs.
@@
Motivated by the potential of GMNs to process diverse NeRF architectures, we investigate whether a GMN encoder can learn a latent space where distances reflect the similarity between the actual content of the radiance fields rather than their specific neural parameterization. Our empirical study reveals that this latent space organization cannot be achieved by solely relying on a rendering loss [38], as such a loss alone causes different NeRF architectures to aggregate into distinct clusters in the embedding space, even when they represent the same underlying object. To overcome this limitation, we draw inspiration from the contrastive learning literature and introduce a SigLIP loss term [39] that places pairs of NeRFs with different architectures representing the same object close to each other in latent space while pushing other pairs further apart. Combined with a rendering loss, this approach enforces an architecture-agnostic embedding space organized by class and instance. Our experimental study features three families of popular NeRF architectures, namely MLPs, tri-planes [8], and multi-resolution hash tables [27], for a total of 13 diverse architectures, and reveals that our encoder produces latent representations that serve as effective inputs for downstream tasks such as classification and retrieval. Notably, our framework is the first to perform tasks on NeRF parameterized as hash tables by processing their weights. Fig. 1 outlines the key components of our framework.
@@
  • We present the first framework to perform tasks on NeRFs parameterized by diverse architectures by processing their weights.
@@
  • We show that, within the families seen at training time, our framework can effectively process unseen NeRF architectures.
  • We achieve superior results to previous methods operating on single architectures.
Figure 1: Framework overview. Our representation learning framework leverages a Graph Meta-Network [23] encoder to map weights of NeRFs with diverse architectures to a latent space where NeRFs representing similar objects are close to each other, regardless of their architecture. The embeddings are then used as input to downstream pipelines for either classification or retrieval tasks.
@@ 2 Related Work
Neural Radiance Fields. NeRFs were first introduced by Mildenhall et al. [26] as a method for novel view synthesis, namely the task of generating previously unseen views of a scene from a set of sparse input images taken from different viewpoints. In the original formulation, a NeRF is an MLP that parameterizes a function $(x,y,z,\theta,\phi)\mapsto(r,g,b,\sigma)$ that maps a 3D position $(x,y,z)$ and a 2D viewing direction $(\theta,\phi)$ to an emitted color $(r,g,b)$ and volume density $\sigma$. Since then, several architectural variants have been proposed, many of which combine the MLP with a trainable discrete data structure that quantizes the space of input coordinates and maps them to a higher-dimensional vector, which becomes the actual MLP input. These structures include voxel grids [24], tri-planes [8], and multi-resolution hash tables [27], and typically result in NeRF architectures that can be trained much faster to convergence without sacrificing rendering quality. In particular, hash tables are the most widely adopted NeRF architecture in recent works [5, 35, 19, 21, 11]. This paper focuses on the NeRF architectures used in Zama Ramirez et al. [38] and Cardace et al. [7], consisting of a single MLP and a tri-plane followed by an MLP, respectively, and the one proposed by Müller et al. [27], i.e. a multi-resolution hash table followed by an MLP. In all three cases, the MLP adheres to the simplified formulation by Mildenhall et al. [26] with no viewing direction, i.e. $(x,y,z)\mapsto(r,g,b,\sigma)$.
@@
Meta-networks. Due to the high dimensionality of the weight space, its symmetries [17], and the impact of randomness on the solution where training converges [14, 1], processing neural network weights presents unique challenges that set them apart from more common input formats. The first works to address the design of neural networks that ingest the weights of other neural networks leverage group theory to devise architectures that are equivariant to the permutation symmetries of the input networks [28, 40, 41]. Yet, these meta-networks are tailored to specific input networks, such as MLPs and CNNs without normalization layers, and cannot generalize to arbitrary input architectures. To overcome this limitation, Graph Meta-Networks (GMNs) were introduced [23, 20]. Since GMNs are graph neural networks, they are, by design, equivariant to the node permutations of input graphs and can ingest any type of graph. Therefore, the challenge of processing neural network weights turns into the task of transforming the input network into a graph. In this paper, we use the GMN formulation by Lim et al. [23] and devise our own conversion of hash tables into graphs.
@@
Meta-networks for NeRF processing. Being the meta-network literature still in its infancy, none of the aforementioned works include NeRFs as input in their experimental evaluation and choose instead to focus on simpler neural networks. The first methods to perform tasks on NeRFs by ingesting their weights are nf2vec [38] and the framework by Cardace et al. [7]. nf2vec is an encoder-decoder architecture trained end-to-end with a rendering loss; at inference time, the encoder takes the weights of a NeRF as input and produces an embedding, which in turn becomes the input to traditional deep learning pipelines for downstream tasks. More recent works [4, 2] investigate the potential applications of this approach to language-related tasks. While nf2vec is designed to ingest MLPs, Cardace et al. processes tri-planar NeRFs [8] by discarding the MLP and giving the tri-planar component alone as input to a Transformer. Yet, both nf2vec and Cardace et al. suffer from the same drawback as the first meta-networks: they are designed to handle specific NeRF architectures. In this work, we instead present the first architecture-agnostic framework for NeRF processing.
@@
Contrastive learning. Contrastive learning is a representation learning approach that trains models to distinguish between similar (positive) and dissimilar (negative) data pairs by aiming to embed similar data points closer together while pushing dissimilar ones farther apart in latent space. [10, 16]. Multimodal vision-language models extend this concept to align image and text modalities by maximizing the similarity between matching image-text pairs and minimizing it for mismatched ones, as demonstrated in CLIP [30], a model that learns a shared embedding space supporting zero-shot transfer to diverse vision tasks. Zhai et al. [39] build upon this foundation and propose to replace the softmax-based loss used in CLIP with a simple pairwise sigmoid loss, called SigLIP, which is shown to work better for relatively small (4k–32k) batch sizes. In this paper, we use the SigLIP loss to align GMN embeddings of different NeRF architectures representing the same object.
@@ 3 Method
Figure 2: Method overview. Left: parameter graph construction for an MLP (left), a tri-plane (middle), and a multi-resolution hash table (right). For better clarity, the graphs of a single $2\times 2\times 2$ plane and of two $4\times 2$ hash tables are shown. Right: our framework leverages a Graph Meta-Network [23] encoder alongside the nf2vec decoder [38] and is trained end-to-end on a dataset of NeRFs with different architectures ($\mathcal{N}_{j}^{A}$, $\mathcal{N}_{j}^{B}$) with both a rendering ($\mathcal{L}_{\text{R}}$) and a contrastive ($\mathcal{L}_{\text{C}}$) loss.
@@
We tackle the challenging problem of embedding NeRFs parameterized by different neural architectures with a representation learning framework. Our method uses an encoder and a decoder trained end-to-end with a combination of rendering and contrastive objectives, where the encoder is implemented as a Graph Meta-Network (GMN). After training, the frozen encoder converts NeRF weights into an embedding that can serve as input to standard deep learning pipelines for downstream tasks. In the remainder of this section, we will describe each framework component; further details are provided in Appendix B.
@@
From NeRFs to graphs. In order for a NeRF to be ingested by the encoder, it must be converted into a graph. The naive approach to perform this conversion would be to adopt the standard computation graph formulation, namely representing a neural network as a Directed Acyclic Graph (DAG), where nodes are activations and edges hold weight values. However, computation graphs scale poorly with the number of activations in networks with weight-sharing schemes, as a single weight requires multiple edges, one for each activation it affects. This limitation has motivated Lim et al. [23] to introduce the parameter graph representation, where each weight is associated with a single edge of the graph instead of multiple edges. Lim et al. describe the parameter graph construction of several common neural layers, which can then be concatenated to form the overall parameter graph. Among these layers, they provide the parameter graph conversion of spatial parameter grids such as those often used in NeRF architectures, specifically tri-planar ones [8]; they do not, however, extend their parameter graph formulation to multi-resolution hash tables [27]. In this paper, we leverage Lim et al.’s graph representations of linear layers and tri-planes, detailed in Appendix A, and propose a conversion into parameter graphs for hash tables, described here. These three graph representations are shown in Fig. 2 (left). Specifically, hash tables map via a hashing function spatial locations of a dense 3D grid to feature vectors stored in a table; in practice, multiple separate hash tables are used to index grids at different resolutions (hence the multi-resolution terminology). To compute the output, the feature vectors of the grid vertices enclosing a query point $(x,y,z)$ are interpolated at each resolution, and the results concatenated to serve as input to an MLP. The most direct way to convert a hash table into a parameter graph would be to explicitly model the 3D grids of features encoded by the tables and adopt a graph representation analogous to the one used for tri-planes. Albeit simple, explicitly modeling the underlying voxel grid would require as many nodes as there are spatial locations, i.e. it would scale cubically with the resolution. Instead, we construct the parameter graph of a hash table with a node for each table entry and a node for each feature vector dimension; then, we connect each entry node to each feature node via an edge storing the corresponding feature value. This subgraph construction is repeated for each hash table, i.e. for each resolution level. Finally, the concatenation of all feature nodes becomes the first layer of the MLP parameter graph that follows. This graph representation preserves the memory efficiency of the hash table and is the one we adopt.
@@
Encoder. Our framework’s encoder is the GMN proposed by Lim et al. [23], i.e. a standard message-passing Graph Neural Network (GNN) [6] with node and edge features but no global graph-level feature. Node features are updated by message-passing along neighbors and contribute to updating the features of the edges connecting them. The final embedding is obtained via an average pooling of the edge features. Notably, as the encoder is a GNN, it can process any input graph and, hence, any neural network that has undergone the previously described parameter-graph conversion, thereby allowing our framework to handle any NeRF architecture for which a graph representation is known.
@@
Decoder. Our framework leverages the decoder first introduced by nf2vec [38], which takes as input the concatenation of the embedding produced by the encoder alongside a frequency encoding [26] of a 3D point $(x,y,z)$ and outputs a learned approximation of the radiance field value $(r,g,b,\sigma)$ at that point. Thus, the combination of decoder and embedding can itself be seen as a conditioned neural radiance field. Inspired by Park et al. [29], the decoder architecture is a simple succession of linear layers intertwined with ReLU activations and a single skip connection from the input to halfway through the network.
@@
Training. The encoder and the decoder are trained end-to-end with a combination of two loss terms: a rendering loss and a contrastive loss. The rendering loss is the one used by Zama Ramirez et al. [38] to train the nf2vec framework and can be described as follows. Consider a NeRF $\mathcal{N}\colon\mathbf{x}\mapsto(\mathbf{c},\sigma)$ with $\mathbf{x}=(x,y,z)$ and $\mathbf{c}=(r,g,b)$ and let $I\in\mathcal{I}$ be one of the images $\mathcal{N}$ was trained on, with corresponding camera pose and intrinsic parameters. Let $\mathcal{X}=\{\mathbf{x}_{i}\}$ be the set of points sampled along a ray cast from image $I$ into the scene and passing through pixel $\mathbf{p}$ on the image plane, and let $C_{\mathcal{N}}(\mathbf{p})$ be the color computed via volume rendering by accumulating the contributions of the $(\mathbf{c}_{i},\sigma_{i})$ output values of $\mathcal{N}$ for all inputs $\mathbf{x}_{i}\in\mathcal{X}$. Analogously, let $C_{\text{D}}(\mathbf{p})$ be the color computed with the output values produced by the decoder when the encoder takes $\mathcal{N}$ as input. The rendering loss associated with NeRF $\mathcal{N}$ is defined as
@@
[…]
@@
where $\mathcal{L}_{\text{R}}(\mathcal{N})$ is defined in Eq. 1. Our contrastive loss follows instead the definition by Zhai et al. [39], namely
@@
where $\ell_{jk}=1$ if $j=k$ and $-1$ otherwise, $t$ and $b$ are learnable scalar hyperparameters, and $\mathbf{u}_{j}$ and $\mathbf{v}_{k}$ are the L2-normalized encoder embeddings of $\mathcal{N}_{j}^{A}$ and $\mathcal{N}_{k}^{B}$, respectively. Finally, the combined loss used to train our framework is
@@
where $\lambda$ is a fixed hyperparameter. The rationale behind the choice of this loss is discussed in Section 4.1. Fig. 2 (right) shows an overview of our training procedure.
Inference. At inference time, a single forward pass of the encoder converts the parameter graph of a NeRF into a latent vector, which we then use as input to standard deep learning pipelines for classification and retrieval tasks, as outlined in Fig. 1 and detailed in Sections 4.2 and 4.3.

### v3

@@ Abstract
Neural Radiance Fields (NeRFs) have emerged as a groundbreaking paradigm for representing 3D objects and scenes by encoding shape and appearance information into the weights of a neural network. Recent studies have demonstrated that these weights can be used as input for frameworks designed to address deep learning tasks; however, such frameworks require NeRFs to adhere to a specific, predefined architecture. In this paper, we introduce the first framework capable of processing NeRFs with diverse architectures and performing inference on architectures unseen at training time. We achieve this by training a Graph Meta-Network within an unsupervised representation learning framework, and show that a contrastive objective is conducive to obtaining an architecture-agnostic latent space. In experiments conducted across 13 NeRF architectures belonging to three families (MLPs, tri-planes, and, for the first time, hash tables), our approach demonstrates robust performance in classification, retrieval, and language tasks involving multiple architectures, even unseen at training time, while also matching or exceeding the results of existing frameworks limited to single architectures.
@@ 1 Introduction
Neural Radiance Fields (NeRFs) (Mildenhall et al., 2020) have emerged over the last few years as a new paradigm for representing 3D objects and scenes (Xie et al., 2022). A NeRF is a neural network trained on a collection of images to map 3D coordinates to color and density values, which can then be used to synthesize novel views of the underlying object or scene via volume rendering. Due to their continuous nature, NeRFs can encode an arbitrary number of images at any resolution into a finite number of neural network weights, thus decoupling the number of observations and their spatial resolution from the memory required to store the 3D representation. As a result, NeRFs hold the potential to become a standard tool for storing and communicating 3D information, as supported by the recent publication of several NeRF datasets (De Luigi et al., 2023; Hu et al., 2023; Zama Ramirez et al., 2024; Cardace et al., 2024; Amaduzzi et al., 2025).
@@
With the rise of NeRFs as a new data format, whether and how it is possible to perform traditional deep learning tasks on them has become an increasingly relevant research question. The naive solution to this problem involves rendering views of the underlying object from its NeRF representation and leveraging existing neural architectures designed to process images. However, this procedure requires additional computation time and several decisions that are likely to impact its outcome, such as the number of views to render, their viewpoint, and their resolution. A more elegant and efficient approach, explored in recent works, relies on performing tasks on NeRFs by processing their weights as input, thereby requiring no rendering step. This strategy is adopted by nf2vec (Zama Ramirez et al., 2024), a representation learning framework that learns to map NeRF weights to latent vectors by minimizing a rendering loss, where NeRFs are standard Multi-Layer Perceptrons (MLPs). These vectors are then used as input to deep learning pipelines for downstream tasks. A related approach is proposed by Cardace et al. (2024), who, instead of employing traditional MLPs, leverage tri-planar NeRFs (Chan et al., 2022), where input coordinates are projected onto three orthogonal planes of learnable features to compute the input for an MLP.
@@
Both nf2vec (Zama Ramirez et al., 2024) and the method by Cardace et al. (2024), however, are designed to ingest a specific type of NeRF architecture (i.e. MLPs with fixed hidden dimensions in nf2vec and tri-planes with fixed spatial resolution in Cardace et al.), and are thus unable to process diverse input architectures. In the context of a research domain where new NeRF designs are constantly being explored (Xie et al., 2022), this assumption strongly limits their applicability. The issue of handling arbitrary architectures has recently been studied in the broader research field on meta-networks, i.e. neural networks that process other neural networks as input. Specifically, Graph Meta-Networks (GMNs) have been proposed (Lim et al., 2024; Kofinas et al., 2024), namely Graph Neural Networks (GNNs) that can ingest any neural architecture as long as it can be first converted into a graph. Yet, these works do not experiment with NeRFs as input to GMNs.
@@
Motivated by the potential of GMNs to process diverse NeRF architectures, we investigate whether a GMN encoder can learn a latent space where distances reflect the similarity between the actual content of the radiance fields rather than their specific neural parameterization. Our empirical study reveals that this latent space organization cannot be achieved by solely relying on a rendering loss (Zama Ramirez et al., 2024), as such a loss alone causes different NeRF architectures to aggregate into distinct clusters in the embedding space, even when they represent the same underlying object. To overcome this limitation, we draw inspiration from the contrastive learning literature and introduce a SigLIP loss term (Zhai et al., 2023) that places pairs of NeRFs with different architectures representing the same object close to each other in latent space while pushing other pairs further apart. Combined with a rendering loss, this approach enforces an architecture-agnostic embedding space organized by class and instance. Our experimental study features three families of popular NeRF architectures, namely MLPs, tri-planes (Chan et al., 2022), and multi-resolution hash tables (Müller et al., 2022), for a total of 13 diverse architectures, and reveals that our encoder produces latent representations that serve as effective inputs for downstream tasks such as classification, retrieval, and language tasks. Notably, our framework is the first to perform tasks on NeRF parameterized as hash tables by processing their weights. Fig.˜1 outlines the key components of our framework.
@@
  • We present the first framework that performs tasks on NeRFs parameterized by diverse architectures by processing their weights.
@@
  • We show that, within the families seen at training time, our framework can effectively process unseen NeRF architectures and generalize to NeRFs trained on unseen data.
  • To achieve these goals, we created and will publicly release several new datasets of NeRFs parameterized by diverse architectures.
  • We achieve comparable or superior results to previous methods operating on single architectures.
Figure 1: Framework overview. Our representation learning framework leverages a Graph Meta-Network (Lim et al., 2024) encoder to map weights of NeRFs with diverse architectures to a latent space where NeRFs representing similar objects are close to each other, regardless of their architecture. The embeddings are then used as input to downstream pipelines for classification, retrieval, and language tasks.
@@ 2 Related Work
Neural Radiance Fields. NeRFs were first introduced by Mildenhall et al. (2020) as a method for novel view synthesis, namely the task of generating previously unseen views of a scene from a set of sparse input images taken from different viewpoints. In the original formulation, a NeRF is an MLP that parameterizes a function $(x,y,z,\theta,\phi)\mapsto(r,g,b,\sigma)$ that maps a 3D position $(x,y,z)$ and a 2D viewing direction $(\theta,\phi)$ to an emitted color $(r,g,b)$ and volume density $\sigma$. Since then, several architectural variants have been proposed, many of which combine the MLP with a trainable discrete data structure that quantizes the space of input coordinates and maps them to a higher-dimensional vector, which then serves as the actual MLP input. These structures include voxel grids (Liu et al., 2020), tri-planes (Chan et al., 2022), and multi-resolution hash tables (Müller et al., 2022), and typically result in NeRF architectures that can be trained much faster to convergence without sacrificing rendering quality. In particular, hash tables are the most widely adopted NeRF architecture in recent works (Barron et al., 2023; Wang et al., 2024; Hu et al., 2024; Lee et al., 2025; Chen et al., 2024). This paper focuses on the NeRF architectures used in Zama Ramirez et al. (2024) and Cardace et al. (2024), consisting of a single MLP and a tri-plane followed by an MLP, respectively, and the one proposed by Müller et al. (2022), i.e. a multi-resolution hash table followed by an MLP. In all three cases, the MLP adheres to the simplified formulation by Mildenhall et al. (2020) with no viewing direction, i.e. $(x,y,z)\mapsto(r,g,b,\sigma)$.
@@
Meta-networks. Due to the high dimensionality of the weight space, its symmetries (Hecht-Nielsen, 1990), and the impact of randomness on the solution where training converges (Entezari et al., 2022; Ainsworth et al., 2023), processing neural network weights presents unique challenges that set them apart from more common input formats. The first works to address the design of neural networks that ingest the weights of other neural networks leverage group theory to devise architectures that are equivariant to the permutation symmetries of the input networks (Navon et al., 2023; Zhou et al., 2023a; b). Yet, these meta-networks are tailored to specific input networks, such as MLPs and CNNs without normalization layers, and cannot generalize to arbitrary input architectures. To overcome this limitation, Graph Meta-Networks (GMNs) were introduced (Lim et al., 2024; Kofinas et al., 2024). Since GMNs are graph neural networks, they are, by design, equivariant to the node permutations of input graphs and can ingest any graph. Therefore, the challenge of processing neural network weights turns into the task of transforming the input network into a graph. In this paper, we use the GMN formulation by Lim et al. (2024) and devise our own conversion of hash tables into graphs.
@@
Meta-networks for NeRF processing. As the meta-network literature is still in its infancy, none of the aforementioned works include NeRFs as input in their experimental evaluation and instead choose to focus on simpler neural networks. The first methods to perform tasks on NeRFs by ingesting their weights are nf2vec (Zama Ramirez et al., 2024) and the framework by Cardace et al. (2024). nf2vec is an encoder-decoder architecture trained end-to-end with a rendering loss; at inference time, the encoder takes the weights of a NeRF as input and produces an embedding, which in turn becomes the input to traditional deep learning pipelines for downstream tasks. More recent works (Ballerini et al., 2024; Amaduzzi et al., 2024) investigate the potential applications of this approach to language-related tasks. While nf2vec is designed to ingest MLPs, Cardace et al. processes tri-planar NeRFs (Chan et al., 2022) by discarding the MLP and giving the tri-planar component alone as input to a Transformer. Yet, both nf2vec and Cardace et al. suffer from the same drawback as the first meta-networks: they are designed to handle specific NeRF architectures. In this work, we instead present the first architecture-agnostic framework for NeRF processing.
@@
Contrastive learning. Contrastive learning is a representation learning approach that trains models to distinguish between similar (positive) and dissimilar (negative) data pairs by aiming to embed similar data points closer together while pushing dissimilar ones farther apart in latent space (Chen et al., 2020; He et al., 2020). Multimodal vision-language models extend this concept to align image and text modalities by maximizing the similarity between matching image-text pairs and minimizing it for mismatched ones, as demonstrated in CLIP (Radford et al., 2021), a model that learns a shared embedding space supporting zero-shot transfer to diverse vision tasks. Zhai et al. (2023) build upon this foundation and propose to replace the softmax-based loss used in CLIP with a simple pairwise sigmoid loss, called SigLIP, which is shown to work better for relatively small (4k–32k) batch sizes. For this reason, in this paper, we use the SigLIP loss to align GMN embeddings of different NeRF architectures representing the same object.
@@ 3 Method
Figure 2: Method overview. Left: parameter graph construction for an MLP (left), a tri-plane (middle), and a multi-resolution hash table (right). For better clarity, the graphs of a single $2\times 2\times 2$ plane and of two $4\times 2$ hash tables are shown. Right: our framework leverages a Graph Meta-Network (Lim et al., 2024) encoder alongside the nf2vec decoder (Zama Ramirez et al., 2024) and is trained end-to-end on a dataset of NeRFs with different architectures ($\mathcal{N}_{j}^{A}$, $\mathcal{N}_{j}^{B}$) with both a rendering ($\mathcal{L}_{\text{R}}$) and a contrastive ($\mathcal{L}_{\text{C}}$) loss.
@@
We tackle the challenging problem of embedding NeRFs parameterized by different neural architectures with a representation learning framework. Our method uses an encoder and a decoder trained end-to-end with a combination of rendering and contrastive objectives, where the encoder is implemented as a Graph Meta-Network (GMN). After training, the frozen encoder converts NeRF weights into an embedding that can serve as input to deep learning pipelines for downstream tasks. In the remainder of this section, we will describe each framework component; further details are provided in App.˜B.
@@
From NeRFs to graphs. In order for a NeRF to be ingested by the encoder, it must be converted into a graph. The naive approach to perform this conversion would be to adopt the standard computation graph formulation, namely representing a neural network as a Directed Acyclic Graph (DAG), where nodes are activations and edges hold weight values. However, computation graphs scale poorly with the number of activations in networks with weight-sharing schemes, as a single weight requires multiple edges, one for each activation it affects. This limitation has motivated Lim et al. (2024) to introduce the parameter graph representation, where each weight is associated with a single edge of the graph, rather than multiple edges. Lim et al. describe the parameter graph construction of several common neural layers, which can then be concatenated to form the overall parameter graph. Among these layers, they provide the parameter graph conversion of spatial parameter grids, such as those often used in NeRF architectures, specifically tri-planar ones (Chan et al., 2022). However, they do not extend their parameter graph formulation to multi-resolution hash tables (Müller et al., 2022). In this paper, we leverage Lim et al.’s graph representations of linear layers and tri-planes, detailed in App.˜A, and propose a conversion into parameter graphs for hash tables, described here. These three graph representations are shown in Fig.˜2 (left). Specifically, hash tables map spatial locations of a dense 3D grid to feature vectors stored in a table via a hashing function; in practice, multiple separate hash tables are used to index grids at different resolutions (hence the multi-resolution terminology). To compute the output, the feature vectors of the grid vertices enclosing a query point $(x,y,z)$ are interpolated at each resolution, and the results are concatenated to serve as input to an MLP. The most direct way to convert a hash table into a parameter graph would be to explicitly model the 3D grids of features encoded by the tables and adopt a graph representation analogous to the one used for tri-planes. Albeit simple, explicitly modeling the underlying voxel grid would require as many nodes as there are spatial locations, i.e. it would scale cubically with the resolution. Instead, we construct the parameter graph of a hash table with a node for each table entry and a node for each feature vector dimension. Then, we connect each entry node to each feature node via an edge that stores the corresponding feature value. This subgraph construction is repeated for each hash table, i.e. for each resolution level. Finally, the concatenation of all feature nodes becomes the first layer of the MLP parameter graph that follows. This graph representation preserves the memory efficiency of the hash table and is the one we adopt.
@@
Encoder. Our framework’s encoder is the GMN proposed by Lim et al. (2024), i.e. a standard message-passing Graph Neural Network (GNN) (Battaglia et al., 2018) with node and edge features but no global graph-level feature. Node features are updated by message-passing along neighbors and contribute to updating the features of the edges connecting them. The final embedding is obtained via an average pooling of the edge features. Notably, as the encoder is a GNN, it can process any input graph and, hence, any neural network that has undergone the previously described parameter-graph conversion, thereby allowing our framework to handle any NeRF architecture for which a graph representation is known.
@@
Decoder. Our framework leverages the decoder first introduced by nf2vec (Zama Ramirez et al., 2024), which takes as input the concatenation of the embedding produced by the encoder alongside a frequency encoding (Mildenhall et al., 2020) of a 3D point $(x,y,z)$ and outputs a learned approximation of the radiance field value $(r,g,b,\sigma)$ at that point. Thus, the combination of decoder and embedding can itself be seen as a conditioned neural radiance field. Inspired by Park et al. (2019), the decoder architecture consists of a simple succession of linear layers, intertwined with ReLU activations, and a single skip connection from the input to halfway through the network.
@@
Training. The encoder and the decoder are trained end-to-end with a combination of two loss terms: a rendering loss and a contrastive loss. The rendering loss is the one used by Zama Ramirez et al. (2024) to train the nf2vec framework and can be described as follows. Consider a NeRF $\mathcal{N}\colon\mathbf{x}\mapsto(\mathbf{c},\sigma)$ with $\mathbf{x}=(x,y,z)$ and $\mathbf{c}=(r,g,b)$ and let $I\in\mathcal{I}$ be one of the images $\mathcal{N}$ was trained on, with corresponding camera pose and intrinsic parameters. Let $\mathcal{X}=\{\mathbf{x}_{i}\}$ be the set of points sampled along a ray cast from image $I$ into the scene and passing through pixel $\mathbf{p}$ on the image plane, and let $C_{\mathcal{N}}(\mathbf{p})$ be the color computed via volume rendering by accumulating the contributions of the $(\mathbf{c}_{i},\sigma_{i})$ output values of $\mathcal{N}$ for all inputs $\mathbf{x}_{i}\in\mathcal{X}$. Analogously, let $C_{\text{D}}(\mathbf{p})$ be the color computed with the output values produced by the decoder when the encoder takes $\mathcal{N}$ as input. The rendering loss associated with NeRF $\mathcal{N}$ is defined as
@@
where $\mathcal{S}(\mathcal{I})$ is a subset of pixels sampled from $\mathcal{I}$ and $\operatorname{smoothL1}$ is the smooth L1 loss (Girshick, 2015). More precisely, Zama Ramirez et al. compute one rendering loss term for foreground and one for background pixels, and express $\mathcal{L}_{\text{R}}$ as a weighted sum of the two. Let $\mathcal{M},\mathcal{T},\mathcal{H}$ be three NeRF architectures and consider a dataset of NeRFs of 3D objects where each object $j$ appears three times: as a NeRF $\mathcal{M}_{j}$ parameterized by $\mathcal{M}$, as a NeRF $\mathcal{T}_{j}$ parameterized by $\mathcal{T}$, and as a NeRF $\mathcal{H}_{j}$ parameterized by $\mathcal{H}$. Given a mini-batch $\mathcal{B}=\{(\mathcal{N}_{1}^{A},\mathcal{N}_{1}^{B}),(\mathcal{N}_{2}^{A},\mathcal{N}_{2}^{B}),\dots\}$, where each $(\mathcal{N}_{j}^{A},\mathcal{N}_{j}^{B})$ is randomly sampled from $\{(\mathcal{M}_{j},\mathcal{T}_{j}),(\mathcal{T}_{j},\mathcal{H}_{j}),(\mathcal{M}_{j},\mathcal{H}_{j})\}$, the rendering loss computed over $\mathcal{B}$ is the average rendering loss of each NeRF $\mathcal{N}$ in $\mathcal{B}$, i.e.
@@
where $\mathcal{L}_{\text{R}}(\mathcal{N})$ is defined in Eq.˜1. Our contrastive loss follows instead the definition by Zhai et al. (2023), namely
@@
where $\ell_{jk}=1$ if $j=k$ and 1-1 otherwise, $t$ and $b$ are learnable scalar hyperparameters, and $\mathbf{u}_{j}$ and $\mathbf{v}_{k}$ are the L2-normalized encoder embeddings of $\mathcal{N}_{j}^{A}$ and $\mathcal{N}_{k}^{B}$, respectively. Finally, the combined loss used to train our framework is
@@
where $\lambda$ is a fixed hyperparameter. The rationale behind the choice of this loss is discussed in Sec.˜4.1. Fig.˜2 (right) shows an overview of our training procedure.
Inference. At inference time, a single forward pass of the encoder converts the parameter graph of a NeRF into a latent vector, which we then use as input to deep learning pipelines for classification, retrieval, and language tasks, as outlined in Fig.˜1 and detailed in Secs.˜4.2, 4.3 and $4.$